OpenStudy (anonymous):

Find the equation of a circle in standard form with center at the point (-3,2) and tangent to the line (touching the line) y=4.

OpenStudy (anonymous):

y=4 is a distance of 2 from (-3, 2), which tells you the radius is 2. Formula for a circle is (x-x1)^2+(y-y1)^2=r^2, or (x+3)^2+(y-2)^2=4